1] A newborn child receives a $20000 gift towards her tertiary education from her grandparents who insisted the gift should be put in investment for 18 years by the baby’s parents. The highest rate offered was 7%. Parents of the baby opted for the investment to be compounded quarterly.


i. What is the compound interest on the investment


ii. How much is the investment worth.



1
Expert's answer
2022-03-21T12:18:13-0400

1.The compound interest on the investment

Compound interest = "=P[(1+i)^n-1]"

where the principal amount (P) = 20000

the nominal interest rate (i) = "\u200b\n \\frac{0.07}{4}"

number of compounding periods (n) "=18*4=72"

Compound interest = "20000[(1+\n\u200b\n \\frac{0.07}{4})"72 -1]

"=20000*2.487209897"

"= 49744.19794"

The compound interest on the investment is approximately $49744.20

2.The investment worth

The investment worth (A) = P + I

where I is the compound interest

A"= 20000+49744.20"

"=69744.20"

The investment is worth approximately $69744.20